OpenStudy (anonymous):

Solve for x: -3|x - 3| = -6

OpenStudy (anonymous):

I would isolate the |x-3| to one side of the equation and then use the properties of absolute values to solve for x.

OpenStudy (anonymous):

I think x=5 because: -3(x-3)=-6 -3x-(-9)=-6 -3x+9=-6 -6-9=-15 -15+9=-6 x=5... I think :)

OpenStudy (anonymous):

5 is indeed one solution, but is it the only solution?